OpenStudy (anonymous):

Verify the identity. 4csc2x = 2csc^2x(tanx) PLEASE HELP ME IF YOU KINDLY WOULD:)

OpenStudy (dangerousjesse):

Express both sides in terms of sine and cosine. Write cosecant as 1/sine and tangent as sine/cosine: \(4 \frac{1}{sin(2 x)} = ^?\frac{sin(x)}{cos(x)} 2 ( \frac{1}{sin(x)} ^2 )\) Simplify the right hand side. \(2 (\frac{1}{sin(x)})^2 (\frac{sin(x)}{cos(x)}) = \frac{2}{cos(x) sin(x)}:\) \(\frac{4}{sin(2 x)} = ^?\frac{2}{cos(x) sin(x)}\) Eliminate the denominators on both sides. Cross multiply: \(4 cos(x) sin(x) = ^?2 sin(2 x)\) Cancel common terms. Divide both sides by 2: \(2 cos(x) sin(x) = ^?sin(2 x)\) Use the double angle identity on \(sin(2 x).\) \(sin(2 x) = 2 sin(x) cos(x):\) \(2 cos(x) sin(x) = ^?2 cos(x) sin(x)\) Come to a conclusion. The left hand side and right hand side are identical:

OpenStudy (dangerousjesse):

Which tells us..?

OpenStudy (dangerousjesse):

I'm sorry, I meant to say \(\cos\) and \(\sin\) not \(cos\) and \(sin\)

OpenStudy (anonymous):

It's fine.

OpenStudy (anonymous):

Tells us that the identity is verified and equal to one another?